OpenStudy (anonymous):

A wire that is 0.20 meters long is moved perpendicularly through a magnetic field of strength 0.45 newtons/amperes·meter at a speed of 10.0 meters/second. What is the emf produced?

OpenStudy (shikamaru11):

so first configure the data given........... l=0.20m B=0.45N/Am V=10m/s and we've to find Emf=???

OpenStudy (shikamaru11):

Emf = v * l * B, when v and B are perpendicular to each other..... so sin\[\sin \theta=1\]

OpenStudy (shikamaru11):

hence put your data and here you go for the answer :)

OpenStudy (michele_laino):

the situation, can be represented by this drawing:  after an infinitesimal time \(dt\), the wire has described the dashed rectangle, whose area is: \(dA=vdtL\), so if we apply the law of magnetic induction we get the requested \(emf\), as below: \[E = \frac{{d\phi }}{{dt}} = \frac{{BdA}}{{dt}} = \frac{{B \cdot v \cdot dt \cdot L}}{{dt}} = BvL\] where, \(d\phi\) is the infinitesimal change of magnetic flux of the field \(B\), and \(L\) is the length of the wire, furthermore \(v\) is its speed

OpenStudy (anonymous):

I don't understand ow to do any of the, that's why I was just asking for the answer.

OpenStudy (michele_laino):

the requested value of the \(emf\), is: \[E = Blv = 0.45 \cdot 0.20 \cdot 10 = ...volts\]

OpenStudy (anonymous):

.9?

OpenStudy (michele_laino):

that's right! we have an \(emf\) of \(0.9\) volts